OpenStudy (anonymous):

Using the following equation: 2 NaOH + H2SO4  2 H2O + Na2SO4 How many grams of sodium sulfate will be formed if you start with 200 grams of sodium hydroxide and you have an excess of sulfuric acid?

OpenStudy (abmon98):

Number of Moles Could be Calculated using the equation Mass(g)/Molar Mass(g/mol) Molar Mass of Sodium hydroxide:(23*1)+(16*1)+(1*1)=40g/mol 200/40= 5moles of sodium hydroxide Ratio of Moles of NaOH to Na2SO4 is 2:1 so 5 moles of NaOH reacts then 5*1/2 of Sodium sulphate are produced.

OpenStudy (abmon98):

Molar Mass of Sodium Sulphate: 142 g/mol 2.5=Mass/142 Mass=142*2.5-355grams

OpenStudy (aaronq):

@Abmon98 it's great you know how to do this, but you should let them work the final answer out - explaining it is good enough.